find the greatest common factor of 16 and 20
Answer by stanbon(48568)

find the greatest common factor of 16 and 20
-------
Convert to prime factor form:
16 = 2^4
20 = 2^2*5
------------------
The greatest common factor is 2^2 = 4
=========================================
Cheers,
Stan H.